OpenStudy (anonymous):

evaluate lim{f(2 + h) - f(2)}/{h}, h->0 where f(x) = -3 - sqrt{x}

OpenStudy (jdoe0001):

\(\bf lim_{h \to 0} \cfrac{f(2+h)-f(2)}{h} \qquad \qquad \qquad f(x) = -3-\sqrt{x} \quad ?\)

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

\(\bf lim_{h \to 0} \cfrac{f(2+h)-f(2)}{h} \qquad \qquad \qquad f(x) = -3-\sqrt{x}\\ \quad \\ \quad \\ f(2) = -3-\sqrt{2} \qquad \qquad f(2+h) = -3-\sqrt{2+h}\\\quad \\ \cfrac{f(2+h)-f(2)}{h} \implies \cfrac{-3-\sqrt{2+h}-(-3-\sqrt{2})}{h}\\\quad \\ \cfrac{\cancel{-3}-\sqrt{2+h}\cancel{+3}+\sqrt{2}}{h} \implies \cfrac{-\sqrt{2+h}+\sqrt{2}}{h} \implies \cfrac{\sqrt{2}-\sqrt{2+h}}{h}\\ \textit{now let's us multiply for the conjugate of the numerator}\\ \cfrac{\sqrt{2}-\sqrt{2+h}}{h} \cdot \cfrac{\sqrt{2}+\sqrt{2+h}}{\sqrt{2}+\sqrt{2+h}} \implies \cfrac{\cancel{2}-(\cancel{2}+\cancel{h})}{\cancel{h}\cdot (\sqrt{2}+\sqrt{2+h})}\)

OpenStudy (anonymous):

so it would be \[1/\sqrt{4}=== 1/2\]

OpenStudy (jdoe0001):

\(\bf \sqrt{2} + \sqrt{2} \implies 2\sqrt{2} \)

OpenStudy (jdoe0001):

\(\bf \sqrt{2} \times \sqrt{2} \implies \sqrt{4} \)

OpenStudy (anonymous):

so when u add it, one of the square symbols cancels?

OpenStudy (jdoe0001):

well, no, they just add, like you'd with a variable say \(\bf a = \sqrt{2}\\ a + a = 2a\)

OpenStudy (anonymous):

i got it thank, before u go i got one more question.

OpenStudy (anonymous):

If f(x) = 19, find f'( 11 )

OpenStudy (anonymous):

i tried using f(x+h)−f(x)/h equation but idk

OpenStudy (jdoe0001):

hmmm... f'(x) = 0 dy/dx of a constant is 0

OpenStudy (anonymous):

thanks for ur help

OpenStudy (jdoe0001):

yw